OpenStudy (anonymous):

Alright I'm confused. Use 1-1 property of the exponential function to solve the given equation. 3^x+2=729/3^x I'm clueless

OpenStudy (loser66):

To me, time both sides by 3^x to get \(3^x(3^x+2)=729\) then plus 1 both sides to get \((3x +1)^2=730\)

OpenStudy (loser66):

sorry, \((3^x+1)^2=730\)

OpenStudy (loser66):

square root both sides to get \(3^x +1 = \sqrt{730}\) you can step up from this, right?

OpenStudy (anonymous):

Yeah I see what you're saying

OpenStudy (anonymous):

We have to solve it using log( )though

OpenStudy (anonymous):

I have no idea.. mlp!

OpenStudy (usukidoll):

oh hey lol

OpenStudy (anonymous):

So I got to log3(3^x-2)=log3(729\3x) But I seem to be lost past that

OpenStudy (loser66):

It seems you need some more steps. -2 both sides, \(3^x= \sqrt{730}-2\) therefore \(\large log_3({\sqrt{730}-2})=x\)

OpenStudy (loser66):

@jdoe0001

OpenStudy (loser66):

help me, friend.

OpenStudy (anonymous):

let me retype this in an easier way to understand lol I just learned the power alt keys and what not

OpenStudy (anonymous):

3ⁿ+²=729/3ⁿ

OpenStudy (loser66):

oh, different problem.

OpenStudy (anonymous):

That plus is part of the power, I just couldn't make it look like it lol

OpenStudy (loser66):

\(\large 3^{n+2}= \dfrac{729}{3^n}\) right?

OpenStudy (loser66):

\(3^{n+2}= 3^n*3^2=3^n *9\) got me?

OpenStudy (anonymous):

There you go!

OpenStudy (loser66):

I am waiting for your reply

OpenStudy (anonymous):

yeah I got it so far

OpenStudy (loser66):

divided both sides by 9 what do you have?

OpenStudy (anonymous):

3^x=81?

OpenStudy (loser66):